OpenStudy (anonymous):

How do I find the regions for the triple integral to calculate the mass of the the volume between the paraboloid x^2+y^2 = 2az and the sphere x^2+y^2+z^2=3a^2 with a density of d(x,y,z)=x^2+y^2+z^2?

OpenStudy (amistre64):

is that similar to finding the solution to a system of equations in order to compute that volume of a solid?

OpenStudy (anonymous):

I didn't quite understand what you mean... To be more clear, I must find out the volume between the two objects, then calculate it's mass. (oh and a>0)

OpenStudy (amistre64):

oh... so the sphere and the paraboloid arent intersecting are they...lol. Wrong concept on my part :)

OpenStudy (anonymous):

yes they are intersecting, and I must find that volume.

OpenStudy (amistre64):

when two curves meet and we want to find the area between them, we integrate the bounds between where they meet; is that something we can do to the sphere and parabaloid?

OpenStudy (amistre64):

when does: x^2 +y^2 -2az = x^2 +y^2 +z^2-3az^2 ?

OpenStudy (anonymous):

it intersects in a circle in the plane z=a

OpenStudy (amistre64):

that circle is one bound of the interecting shapes then; do we have an x bound and a ybound yet?

OpenStudy (amistre64):

we are trying, if i see it correctly, find the area that is scooped out of an ice cream container :)

OpenStudy (amistre64):

area means volume lol

OpenStudy (amistre64):

there should be an upper and lower bound of all three dimensions

OpenStudy (amistre64):

if we could extablish a cross section; could we spin it to find the volume of the solid created by the rotation about an axis?

OpenStudy (anonymous):

Unusually complicated. I would try\[\int\limits_{0}^{2\pi}\int\limits_{r ^{2}/2a}^{\sqrt{3a ^{2}-r ^{2}}}\int\limits_{a \sqrt{2(3a ^{2}-1)}}^{a ^{2\sqrt{6}}}r ^{3+z ^{2}drdzd \theta}\]

OpenStudy (anonymous):

thanks guys, I'll give it a try. Although I'm not so sure about the limits of dr